OpenStudy (anonymous):

Graph the equation 3x+y=1 using the slope and y-intercept.

OpenStudy (saifoo.khan):

y = -3x + 1 Slope = -3 y intercpt = 1

OpenStudy (anonymous):

y=-3x+1 or written as y=-(3/1)x+1 y=mx+b m= (slope) rise over run b=the point of y intercept